Find the ration of de Broglie wavelength of molecules of hydrogen and helium which are at temperatures 27°C and 127°C respectively.

Grade:11

1 Answers

Norman Manchu
6 years ago
\frac{\lambda1}{\lambda2}=\frac{h}{m1v1}\frac{m2v2}{h}=\frac{4v2}{2v1}=\frac{4\sqrt{400}\sqrt{2}}{2\sqrt{300}\sqrt{4}}=\sqrt\frac{8}{3}
 
I used formula:v=√3RT|√M
Convert temperature into kelvins K=°C+273
m signifies mass and M signifies molar mass and both are taken equal respective to the molecule since their is only one molecule each as mentioned in the question|